7/v=9/v-1/3 what is the solution for v
Answer by mananth(16946)

7/v=9/v-1/3 what is the solution for v
Add -9/v to both sides
7/v-9/v=9/v-9/v-1/3
-2/v = -1/3
-2*-3 = v*1
v=6